Find the square roots of the following numbers by the Prime Factorisation Method.  viii. 9216  

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

9216 = 2×2×2×2×2×2×2×2×2×2×3×3×1⇒9216=(2×2)×(2×2)×(2×2)×(2×2)×(2×2)×(3×3) ⇒9216=(2×2×2×2×2×3)×(2×2×2×2×2×3) ⇒9216=96×96

⇒9216=(96)2
⇒√9216=96

Thanks